OpenStudy (anonymous):

According to a national survey, 92% of U.S. Dog owners consider their pets to be members of the family. In a random survey of 150 dog owners, what is the probability that at most 132 of them consider their pets to be members of the family?

OpenStudy (anonymous):

You will need to get a second opinion on this... 92% consider their dogs to be members of family 92% of 150 = 0.92 * 150 = 138 132/138 = 0.956 = 95.6% or if rounded = 96%

OpenStudy (anonymous):

but like I said...you will probably need to get a second opinion because I am not the best at probabilities

OpenStudy (anonymous):

That helped thank you! so much! i really appreciate it! @kelliegirl33

OpenStudy (anonymous):

@collinfootball